1) Prove that a matrix M2,3 is a vector space such that A= i_1211 212 213] ^ [a21 222 223]
Answer #1

1 D het, Man be the set of all 2x3 matrices. Here the zero matrix [oo o] EM2, 3 . So, M2, 3 & 4 . het, A, B, C & Maya and x,A + B = autben a 12+ biz Lasit bei Azat bez G3tbi3] az 3 + b23] - bu tau bez tous les tous] . [baitan bertans best aes] = B+A7) a r & (A+B) = . autbil Lazıt bal air+biz azather Cizrbi3 / A3+ b23] = Tacoutbu) a lantbis) La 1821 + be) alaset bie) or (a